OpenStudy (anonymous):

simplifying the rational expression. State any excluded value 3x-6/x-2. I think it's A. 3; where x =2 (the equal has a line going through it) B. x C.6 where x =3 (also has a line going through the equal) One D.1 If i'm wrong let me know please explain the answer ^ ^

OpenStudy (usukidoll):

this? \[\frac{3x-6}{x-2}\]

OpenStudy (usukidoll):

I hope factoring further isn't required because I can cancel out the x-2 if I take the three out so for this particular fraction we need to find an x that makes it undefined, so we need to exclude it . It's like solving x-2=0

OpenStudy (anonymous):

Yeah It looks like that

OpenStudy (anonymous):

Okay, Can you help me solve it, i'm not good with math?

OpenStudy (usukidoll):

for x -2 = 0?

OpenStudy (usukidoll):

this feels like a double edge sword problem... if I factor further the denominator is wiped out

OpenStudy (usukidoll):

if I don't factor, then we can find an x value that makes the fraction undefined.

OpenStudy (anonymous):

I don't have that option >~<

OpenStudy (usukidoll):

ahhhhhhh I found the problem :P one part is what makes that fraction undefined the other is to simplify XD

OpenStudy (usukidoll):

ok a two parter :P what makes x undefined? solve x-2 = 0

OpenStudy (anonymous):

Not really sure 1?

OpenStudy (usukidoll):

x-2= 0 what do we need to do to have x by itself add, subtract, multiply, or divide

OpenStudy (anonymous):

Get rid of add two to both sides?

OpenStudy (usukidoll):

yes so x =2 ... that's our restriction

OpenStudy (usukidoll):

\[\frac{3x-6}{x-2} \] now we just have to factor on the numerator . what number do they have in common? hint: it's the three times table 3 x 1 3 x 2

OpenStudy (anonymous):

3?

OpenStudy (usukidoll):

yeah so we take the 3 out \[\frac{3(x-2)}{x-2}\]

OpenStudy (usukidoll):

then we can cancel out x-2

OpenStudy (anonymous):

That leaves us with 3x?

OpenStudy (usukidoll):

close... just 3

OpenStudy (usukidoll):



OpenStudy (usukidoll):